# InsightBuddy-AI: Medication Extraction and Entity Linking using Large Language Models and Ensemble Learning

### v1

@@ 4.1 Individual Models Finetuned: sub-word level
4.2 Ensemble: word-level grouping max/first/avg-logit
We tried first logit voting, max voting, and average voting to group sub-words into words with corresponding labels. Their results are shown in Figure 12, in the upper cluster. First logit voting produced a higher Recall 0.8260 while Max logit voting produced a higher Precision 0.8261 resulting in higher F1 0.8232, i.e. Max logit > First logit > Average logit with macro F1 (0.8232, 0.8229, 0.8227). However, overall, their performance scores are very close, so we chose the first-logit voting output for the afterward word-level ensemble due to computational convenience.
4.3 Word-level vs Sub-word Level scores
From word-level ensemble result in Figure 13, it says that the ensembled model can achieve word-level evaluation scores 0.826, 0.826, and 0.823 for macro P/R/F1, which is close to sub-word level best model 0.847 F1. We can see that at word-level evaluation, there are 563,329 support tokens in Figure 13, vs sub-word level 756,014 tokens in Figure 8. Can we also report word-level eval without distinguishing B/I? i.e. as long as the label catergies are correct regardless it is B or I.
Word-level ensemble voting, max-logit voting > first-logit > average-logit, as shown in Figure 12, with Macro F1 scores (0.8232, 0.8229, 0.8227) respectively, which are very close though. They have the same weighted average F1 and Accuracy scores (0.9798, 0.9796) respectively.
Figure 12: word-level grouping ensemble voting, max > first > average logit voting though they are very close scores. The stacked ensemble has the highest Precision scores, but the lowest Recall scores, which lead to lower F1. in the bottom cluster, it is the word-level evaltion without distinguishing B/I labels, Eval on n2c2 2018 test data.
@@ 4.3 Word-level vs Sub-word Level scores
4.4 Ensemble: Voting vs Stacked
We use the Max logit stacked ensemble as an example, in figure 15, which shows that the Stacked Ensemble produced much lower evaluation scores macro avg (0.6863 0.7339 0.6592) than the voting mechanism macro avg (0.8261 0.8259 0.8232) for (P, R, F1). The corresponding confusion matrix from the stacked ensemble using the max logic is shown in Figure 16 with more errors spread in the image, the colored numbers outside the diagonal line.
Figure 15: word-level grouping ensemble, max logit: stacked ensemble Eval on n2c2 2018 test data, which is much lower than the max voting.
@@ 4.4 Ensemble: Voting vs Stacked
4.5 Word-level: ensembles vs individual fine-tuned
BioMedRoBERTa individual word level max logit grouping scores macro avg P/R/F1 (0.8065 0.8224 0.8122 563329) vs max logit ensemble voting P/R/F1 (0.8261 0.8259 0.8232), we can see that ensemble boosted P (0.8261-0.8065)/0.8065= 2.43%, and F1 (0.8232-0.8122)/0.8122= 1.35% which says the ensemble voting is successful. By increasing the Precision score, the ensembles reduce the false positive labels in the system output, while keeping the Recall at the same level, i.e. the true positive labels.
Figure 17: word-level individual model (grouping using first-logit) vs ensemble using max-logit, Eval on n2c2 2018 test data
4.6 Ensemble Models: BIO-span vs non-strict word-level
So far, we have been reporting the evaluation scores on the BIO-strict label categorization, i.e. we distinguish between the label’s beginning or the inner part of the label. For instance, a B-Drug will be different from an I-Drug and it will be marked as wrong if they are different from the reference. However, we think in practice, there are situations when users do not need the BIO, especially B and I difference. In Figure 12, we can see that, without considering the label difference of B and I, only focusing on the 9 label categories, word level ensemble model produced much higher Macro avg evaluations cores on Precision (0.8844) and Recall (0.8830) leading to higher F1 (o.8821), in comparison to BI-distinguished Macro F1 0.8232 (voting-max-logit) and F1 0.8156 (stacked-first-logit).
5 Entity Linking: BNF and SNOMED
To map the identified named entities into the clinical knowledge base. We use the existing code mapping sheet from the British National Formulary (BNF) web between, SNOMED-CT, BNF, dm+d, and ICD  [^2] https://www.nhsbsa.nhs.uk/prescription-data/understanding-our-data/bnf-snomed-mapping. We preprocessed the SNOMED code from 377,834 to 10,804 to filter repeated examples between the mapping of SNOMED and BNF. We looked for non-drug words present in the text, then we filtered the drugs further by seeing if words like [’system’, ’ostomy’, ’bag’, ’filter’, ’piece’, ’closure’] were present in the text and if so, it was discarded.
For SNOMED CT mapping, we applied fuzzy search on the cleaned mapping list with drug names. Then we SNOMED CT code will be added to the searching function on the SNOMED CT web, whenever there is a match.
For BNF mapping, the linking function uses keyword search to retrieve the BNF website with corresponding drugs, due to its different searching features in comparison to the SNOMED-CT web page. Users can select whichever is suitable to their preferences between the two clinical knowledge bases (KBs), Figure 18.
Figure 19: Choice of BNF and SNOMED-CT Linking
Figure 20: Demonstration of Clinical Events Outputs using A Synthetic Letter
6 MedNER Desktop Application and Interface: Insight-Buddy-AI
We offer the desktop application of individually fine-tuned models that we developed during this investigation.
Figure 19 shows the options to choose either BNF or SNOMED-CT mapping for Medical Coding usage. Figure 20 demonstrates the NER outcomes using a synthetic clinical letter generated by Claude. Figure 21 shows the interface can load any Huggingface NER models and example output from Model https://huggingface.co/Xenova/bert-base-NER 333we need this conversion script to make the models into .onnx but we will make our own automatic version soon for huggingface https://github.com/xenova/transformers.js/blob/main/scripts/convert.py. Users can do inference on any NER model but they do need it to be .onnx in its folder structure. Figure 19 shows the features of InsightBuddy-AI with context-awareness using a window parameter around the recognised entity.
We also recorded live demos on the desktop applications 444longer video demo https://youtu.be/4NKpcHdxeko and 2min live demo https://youtu.be/rLBHA35jhwk.
The applications currently support file types: PDF, DOCX, DOC, and TXT; operating systems: Both windows and Mac.
7 Discussion and Conclusion
In this paper, we investigated Stacked Ensemble and Voting Ensemble on medical named entity recognition tasks using eight pretrained LMs from both general and biomed/clinical domains. Our experiments show that our fine-tuned best individual models outperformed the state-of-the-art on standard shared task data n2c2-2018. The ensemble strategies further improved the individual model performances. We also offered desktop applications and user interfaces for individual fine-tuned models where we added the entity linking/normalisation function to BNF and SNOMED CT clinical knowledge base. We call the interface InsightBuddy-AI, which is publically available at .
Ethics
Acknowledgments
We thank Sam and Nicolo for the advice and help on this project, espacially their scientific feedback. PR is supported by the Uni Manchester urenco fund; LH and GN are grateful for the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC).
Appendix A Hyper Parameter Optimisations
@@
We list the model evaluation scores using different parameter set on validation data.
Appendix B Individual Model Training Parameters - used at Stage/Version-1

### v2

@@ B.2 Ensemble: Stacked using output logits (non one-hot)
Figure 17: word-level grouping ensemble, max logit (logits, non-one-hot): stacked ensemble Eval on n2c2 2018 test data, which is much lower than the max voting.
@@
B.3 Individual vs Ensemble Models
The word-level performance comparisons from individual models and voting max-logit ensembles are presented in Table LABEL:tab:individual-vs-ensemble-max-voting-word-level.
Appendix C Huggingface Models used
InsightBuddy integrated individual models and their Huggingface repositories are listed in Table LABEL:tab:EnsembleNER-model-list.
Appendix D Hyper Parameter Optimisations
@@
Table 2: word-level individual model (grouping using max-logit) vs ensemble using max-logit, Eval on n2c2 2018 test data
Individual models max-logit grouping (word)
Metric P R F1
BERT
accuracy 0.9773
macro avg 0.7942 0.7965 0.7928
weighted avg 0.9784 0.9773 0.9775
RoBERTa
accuracy 0.9780
macro avg 0.8029 0.8201 0.8094
weighted avg 0.9795 0.9780 0.9784
RoBERTa-Large
accuracy 0.9788
macro avg 0.8091 0.8351 0.8202
weighted avg 0.9802 0.9788 0.9792
ClinicalBERT
accuracy 0.9780
macro avg 0.8087 0.7916 0.7964
weighted avg 0.9785 0.9780 0.9779
BioBERT
accuracy 0.9776
macro avg 0.7972 0.8131 0.8027
weighted avg 0.9787 0.9776 0.9779
BioClinicalBERT
accuracy 0.9776
macro avg 0.7999 0.8090 0.8017
weighted avg 0.9788 0.9776 0.9779
BioMedRoBERTa
accuracy 0.9783
macro avg 0.8065 0.8224 0.8122
weighted avg 0.9797 0.9783 0.9786
PubMedBERT
accuracy 0.9784
macro avg 0.8087 0.8292 0.8166
weighted avg 0.9800 0.9784 0.9788
Voting Max logit ensemble word level
accuracy 0.9796
macro avg 0.8261 0.8259 0.8232
weighted avg 0.9807 0.9796 0.9798
Model Macro P Macro R Macro F Accuracy Tokens(sub-words)
BERT 0.8336 0.8264 0.8283 0.9748 756798
ROBERTa 0.8423 0.8471 0.8434 0.9770 756014
ROBERTa-L 0.8489 0.8606 0.9782 0.8538 756014
PubMedBERT 0.8324 0.8381 0.8339 0.9783 681211
ClinicalBERT 0.8482 0.8245 0.8341 0.9753 796313
BioMedRoBERTa 0.8482 0.8477 0.8468 0.9775 756014
BioClinicalBERT 0.8440 0.8405 0.8406 0.9751 791743
BioBERT 0.8365 0.8444 0.8393 0.9750 791743
Table 3: InsightBuddy individual sub-word level model eval on n2c2-2018 test set. The first group: normal domain PLM; The second group: biomedical PLM. The different numbers of Support are due to the different tokenizers they used – ROBERTa and ROBERTa-L use the same tokenizers, BioClinicalBERT and BioBERT use the same tokenizers, and other models all use different tokenizers; PubMedBERT generated the least number of sub-words/tokens 681,211 while ClinicalBERT generated the largest number of tokens 796,313.
Table 4: InsightBuddy integrated individual models and their Huggingface repositories.
Ensemble List Link
BERT https://huggingface.co/google-bert/bert-base-uncased
BioBERT https://huggingface.co/dmis-lab/biobert-base-cased-v1.2
ClinicalBERT https://huggingface.co/medicalai/ClinicalBERT
BioClinicalBERT https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT
PubMedBERT https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext
BioMedRoBERTa https://huggingface.co/allenai/biomed_roberta_base
RoBERTa https://huggingface.co/FacebookAI/roberta-base
RoBERTa Large https://huggingface.co/FacebookAI/roberta-large